# Erwin: A Tree-based Hierarchical Transformer for Large-scale Physical Systems

### v1

@@ Abstract
Large-scale physical systems defined on irregular grids pose significant scalability challenges for deep learning methods, especially in the presence of long-range interactions and multi-scale coupling. Traditional approaches that compute all pairwise interactions, such as attention, become computationally prohibitive as they scale quadratically with the number of nodes. We present Erwin, a hierarchical transformer inspired by methods from computational many-body physics, which combines the efficiency of tree-based algorithms with the expressivity of attention mechanisms. Erwin employs ball tree partitioning to organize computation, which enables linear-time attention by processing nodes in parallel within local neighborhoods of fixed size. Through progressive coarsening and refinement of the ball tree structure, complemented by a novel cross-ball interaction mechanism, it captures both fine-grained local details and global features. We demonstrate Erwin’s effectiveness across multiple domains, including cosmology, molecular dynamics, and particle fluid dynamics, where it consistently outperforms baseline methods both in accuracy and computational efficiency.
@@ 1 Introduction
Scientific deep learning is tackling increasingly computationally intensive tasks, following the trajectory of computer vision and natural language processing. Applications range from molecular dynamics (MD) (Arts et al., 2023), computational particle mechanics (Alkin et al., 2024b), to weather forecasting (Bodnar et al., 2024), where simulations often involve data defined on irregular grids with thousands to millions of nodes, depending on the required resolution and complexity of the system.
@@
Such large-scale systems pose a significant challenge to existing methods that were developed and validated at smaller scales. For example, in computational chemistry, models are typically trained on molecules with tens of atoms (Kovács et al., 2023), while molecular dynamics simulations can involve well beyond thousands of atoms. This scale disparity might result in prohibitive runtimes that render models inapplicable in high-throughput scenarios such as protein design (Watson et al., 2023) or screening (Fu et al., 2022).
@@
A key challenge in scaling to larger system sizes is that computational methods which work well at small scales break down at larger scales. For small systems, all pairwise interactions can be computed explicitly, allowing deep learning models to focus on properties like equivariance (Cohen & Welling, 2016). However, this brute-force approach becomes intractable as the system size grows. At larger scales, approximations are required to efficiently capture both long-range effects from slowly decaying potentials or multi-scale coupling (Majumdar et al., 2020). As a result, models validated only on small systems often lack the architectural components necessary for efficient scaling.
@@
This problem has been extensively studied in computational many-body physics (Hockney & Eastwood, 2021), where the need for evaluating long-range potentials for large-scale particle systems led to the development of sub-quadratic tree-based algorithms (Barnes & Hut, 1986; Carrier et al., 1988). These methods are based on the intuition that distant particles can be approximated through their mean field effect rather than individual interactions (Pfalzner & Gibbon, 1996). The computation then is structured using hierarchical trees to efficiently organize the computation at multiple scales. While highly popular for numerical simulations, these tree-based methods have seen limited adoption in deep learning due to poor synergy with GPU architectures.
@@
Transformers (Vaswani et al., 2017), on the other hand, employ the highly optimized attention mechanism, which comes with the quadratic cost of computing all-to-all interactions. In this work, we combine the efficiency of hierarchical tree methods with the expressivity of attention to create a scalable architecture for the processing of large-scale particle systems. Our approach leverages ball trees to organize computation at multiple scales, enabling both local accuracy and global feature capture while maintaining linear complexity in the number of nodes.
@@
    • Model expressivity on large-scale multi-scale phenomena (turbulent fluid dynamics)
@@ 2 Related Works: sub-quadratic attention
One way to avoid the quadratic cost of self-attention is to linearize attention by performing it on non-overlapping patches. For data on regular grids, like images, the SwinTransformer (Liu et al., 2021) achieves this by limiting attention to local windows with cross-window connection enabled by shifting the windows. However, for irregular data such as point clouds or non-uniform meshes, one first needs to induce a structure that will allow for patching. Several approaches (Liu et al., 2023; Sun et al., 2022) transform point clouds into sequences, most notably, PointTransformer v3 (Wu et al., 2024), which projects points into voxels and orders them using space-filling curves (e.g., Hilbert curve). While scalable, these curves introduce artificial discontinuities that can break local spatial relationships.
@@
Rather than using a hierarchical decomposition, another line of work proposes cluster attention (Janny et al., 2023; Alkin et al., 2024a). These methods first group points into clusters and aggregate their features at the cluster centroids through message passing or cross-attention. After computing attention between the centroids, the updated features are then distributed back to the original points. While these approaches achieve the quadratic cost only in the number of clusters, they introduce an information bottleneck at the clustering step that may sacrifice fine-grained details and fail to capture features at multiple scales - a limitation our hierarchical approach aims to overcome.
@@ 3.2 Ball tree
To enable efficient implementation, we want to work with perfect binary trees, i.e. trees where all internal nodes have exactly two children and all leaf nodes appear at the same depth. To achieve this, we pad the leaf level of a ball tree with virtual nodes, yielding the total number of nodes $2^{m}$, where $m=\text{ceil}(\log_{2}(n))$.
@@ 3.2.1 Ball tree properties
The contiguous storage property, combined with the fixed size of balls at each level, enables efficient implementation through tensor operations. Specifically, accessing any ball $B\in L_{i}$ simply requires selecting a contiguous sequence of $2^{i}$ indices. For instance, in the example above, for $i=2$, we select a:d and e:h to access the balls. Since the balls are equal, we can simply reshape $L_{0}$ to access any level. This representation makes it particularly efficient to implement our framework’s core operations - ball attention and coarsening/refinement - which we will introduce next.
Another important property of ball trees is that while they cover the whole point set, they are not required to partition the entire space. Coupled with completeness, it means that at each tree level, the nodes are essentially associated with the same scale. This contrasts with other structures such as oct-trees that cover the entire space and whose nodes at the same level can be associated with regions of different sizes:
@@
Figure 2: Ball tree vs Oct-tree construction. Colors highlight the difference in scales for nodes including the same number of points.
@@ 4.1 Ball tree attention
For each ball attention operator, we specify a level $k$ of the ball tree where each ball $B\in L_{k}$ contains $2^{k}$ leaf nodes. The choice of $k$ presents a trade-off: larger balls capture longer-range dependencies, while smaller balls are more resource-efficient. For each ball $B\in L_{k}$, we collect the leaf nodes within B
@@
along with their features from $X_{\mathrm{leaf}}$
@@
As attention is computed independently for each ball $B\in L_{k}$, the computational cost is reduced from quadratic to linear. Precisely, for ball attention, the complexity is $\mathcal{O}(|B|^{2}\cdot\frac{n}{|B|})$, i.e. quadratic in the ball size and linear in the number of balls:
@@
with a learnable parameter $\sigma\in\mathbb{R}$ (Wessels et al., 2024). The term decays rapidly as the distance between two nodes increases which enforces locality and helps to mitigate potential artifacts from the tree building, particularly in cases where distant points are grouped together.
@@
To increase the receptive field of our attention operator, we implement cross-ball connections inspired by the shifted window approach in Swin Transformer (Liu et al., 2021). There, patches are displaced diagonally by half their size to obtain two different image paTreertitioning configurations. This operation can be equivalently interpreted as keeping the patches fixed while sliding the image itself.
Following this interpretation, we rotate the point cloud and construct the second ball tree $T_{\mathrm{rot}}=\{L_{0}^{\mathrm{rot}},...,L_{m}^{\mathrm{rot}}\}$ which induces a permutation $\pi^{\mathrm{rot}}$ of leaf nodes (see Fig. 3, center). We can then compute ball attention on the rotated configuration by first permuting the features according to $\pi^{\mathrm{rot}}$, applying attention, and then permuting back:
@@
The inverse operation, refinement, allocates information from a coarse representation back to finer scales. More precisely, for a ball $B\in L_{k}$, its features are distributed back to the nodes at level $L_{k-l}$ contained within $B$ as
@@ 4.2 Model architecture
We are now ready to describe the details of the main model to which we refer as Erwin [^4] We pay homage to Swin Transformer as our model based on rotating windows instead of sliding, hence Rwin $\rightarrow$ Erwin. (see Fig. 3) - a hierarchical transformer operating on ball trees.
@@
At the embedding phase, we first construct a ball tree on top of the input point cloud and pad the leaf layer to complete the tree, as described in Section 3.2. To capture local geometric features, we employ a small-scale MPNN, which is conceptually similar to PointTransformer’s embedding module using sparse convolution. When input connectivity is not provided (e.g. mesh), we utilize the ball tree structure for a fast nearest neighbor search.
@@
Following a UNet structure (Ronneberger et al., 2015; Wu et al., 2024), Erwin processes features at multiple scales through encoder and decoder paths (Fig. 3, right). The encoder progressively coarsens the ball tree while increasing feature dimensionality to maintain expressivity. The coarsening factor is a hyperparameter that takes values that are powers of $2$. At the decoder stage, the representation is refined back to the original resolution, with skip connections from corresponding encoder levels enabling multi-scale feature integration.
@@ 5 Experiments
Implementation details for all experiments are given in Appendix C. The code is available at anonymized link. Extended experiments are given in Appendix B, including an additional experiment on airflow pressure modelling.
@@
To experimentally evaluate Erwin’s scaling, we learn the power-law [^5] We only use data for $n\geq 1024$ to exclude overhead costs. form $\mathrm{Runtime}=C\cdot n^{\beta}$ by first applying the logarithm transform to both sides and then using the least square method to evaluate $\beta$. The result is an approximately linear scaling with $\beta=1.054$ with $R^{2}=0.999$, see Fig. 5, left. Ball tree construction accounts for only a fraction of the overall time, proving the efficiency of our method for linearizing attention for point clouds.
@@
One of the theoretical properties of our model is that with sufficiently many layers, its receptive field is global. To verify this claim experimentally, for an arbitrary target node, we run the forward pass of Erwin and MPNN and compute gradients of the node output with respect to all input nodes’ features. If the gradient is non-zero, the node is considered to be in the receptive field of the target node. The visualization is provided in Fig. 5, right, where we compare the receptive field of our model with that of MPNN. As expected, the MPNN has a limited receptive field, as it cannot exceed $N$ hops, where $N$ is the number of message-passing layers. Conversely, Erwin implicitly computes all-to-all interactions, enabling it to capture long-range interactions in data.
@@ 5.1 Cosmological simulations
To demonstrate our model’s ability to capture long-range interactions, we use the cosmology benchmark (Balla et al., 2024) which consists of large-scale point clouds representing potential galaxy distributions.
@@
The dataset is derived from N-body simulations that evolve dark matter particles from the early universe to the present time. After the simulation, gravitationally bound structures (halos) are indicated, from which the $5000$ heaviest ones are selected as potential galaxy locations. The halos form local clusters through gravity while maintaining long-range correlations that originated from interactions in the early universe before cosmic expansion, reflecting the initial conditions of simulations.
@@
The input is a point cloud $\mathbf{X}\in\mathbb{R}^{5000\times 3}$, where each row corresponds to a galaxy and column to $x,y,z$ coordinate respectively. The task is a regression problem to predict the velocity of every galaxy $\mathbf{Y}\in\mathbb{R}^{5000\times 3}$. We vary the size of the training dataset from $64$ to $8192$, while the validation and test datasets have a fixed size of $512$. The models are trained using mean squared error loss
$\mathcal{L}=\text{MSE}(\hat{Y},Y)\vspace{-2pt}$
between predicted and ground truth velocities.
@@ Results
The results are shown in Fig. 6. We compare against multiple equivariant (NequIP (Batzner et al., 2021), SEGNN (Brandstetter et al., 2022)) and non-equivariant (MPNN (Gilmer et al., 2017), PointTransformer v3 (Wu et al., 2024)) baselines. In the small data regime, graph-based equivariant models are preferable. However, as the training set size increases, their performance plateaus. We note that this is also the case for non-equivariant MPNN, suggesting that the issue might arise from failing to capture medium to large-scale interactions, where increased local expressivity of the model has minimal impact. Conversely, transformer-based models scale favorably with the training set size and eventually surpass graph-based models, highlighting their ability to capture both small and large-scale interactions. Our model demonstrates particularly strong performance and significantly outperforms other baselines for larger training set sizes.

### v2

@@ Abstract
Large-scale physical systems defined on irregular grids pose significant scalability challenges for deep learning methods, especially in the presence of long-range interactions and multi-scale coupling. Traditional approaches that compute all pairwise interactions, such as attention, become computationally prohibitive as they scale quadratically with the number of nodes. We present Erwin, a hierarchical transformer inspired by methods from computational many-body physics, which combines the efficiency of tree-based algorithms with the expressivity of attention mechanisms. Erwin employs ball tree partitioning to organize computation, which enables linear-time attention by processing nodes in parallel within local neighborhoods of fixed size. Through progressive coarsening and refinement of the ball tree structure, complemented by a novel cross-ball interaction mechanism, it captures both fine-grained local details and global features. We demonstrate Erwin’s effectiveness across multiple domains, including cosmology, molecular dynamics, PDE solving, and particle fluid dynamics, where it consistently outperforms baseline methods both in accuracy and computational efficiency.
@@ 1 Introduction
Scientific deep learning is tackling increasingly computationally intensive tasks, following the trajectory of computer vision and natural language processing. Applications range from molecular dynamics (MD) (Arts et al., 2023) and computational particle mechanics (Alkin et al., 2024b) to weather forecasting (Bodnar et al., 2024), where simulations often involve data defined on irregular grids with thousands to millions of nodes, depending on the required resolution and complexity of the system.
@@
Such large-scale systems pose a significant challenge to existing methods that were developed and validated at smaller scales. For example, in computational chemistry, models are typically trained on molecules with tens of atoms (Kovács et al., 2023), while molecular dynamics simulations often exceed thousands of atoms. This scale disparity might result in prohibitive runtimes that render models inapplicable in high-throughput scenarios such as protein design (Watson et al., 2023) or screening (Fu et al., 2022).
@@
A key challenge in scaling to larger system sizes is that computational methods that work well at small scales break down at larger scales. For small systems, all pairwise interactions can be computed explicitly, allowing deep learning models to focus on properties like equivariance (Cohen & Welling, 2016). However, this brute-force approach becomes intractable as the system size grows. At larger scales, approximations are required to efficiently capture both long-range effects from slowly decaying potentials and multi-scale coupling (Majumdar et al., 2020). As a result, models validated only on small systems often lack the architectural components necessary for efficient scaling.
@@
This problem has been extensively studied in computational many-body physics (Hockney & Eastwood, 2021), where the need for evaluating long-range potentials for large-scale particle systems led to the development of sub-quadratic tree-based algorithms (Barnes & Hut, 1986; Carrier et al., 1988). These methods are based on the intuition that distant particles can be approximated through their mean-field effect rather than individual interactions (Pfalzner & Gibbon, 1996). The computation is then structured using hierarchical trees to efficiently organize operations at multiple scales. While highly popular for numerical simulations, these tree-based methods have seen limited adoption in deep learning due to poor synergy with GPU architectures.
@@
Transformers (Vaswani et al., 2017), on the other hand, employ the highly optimized attention mechanism, which comes with the quadratic cost of computing all-to-all interactions. In this work, we combine the efficiency of hierarchical tree methods with the expressivity of attention to create a scalable architecture for processing large-scale particle systems. Our approach leverages ball trees to organize computation at multiple scales, enabling both local accuracy and global feature capture while maintaining linear complexity in the number of nodes.
@@
    • Expressivity on large-scale phenomena (PDE benchmarks, turbulent fluid dynamics)
@@ 2 Related Works: sub-quadratic attention
One way to avoid the quadratic cost of self-attention is to linearize attention by performing it on non-overlapping patches. For data on regular grids, like images, the Swin Transformer (Liu et al., 2021) achieves this by limiting attention to local windows with cross-window connections enabled by shifting the windows. However, for irregular data such as point clouds or non-uniform meshes, one first needs to induce a structure that will allow for patching. Several approaches (Liu et al., 2023; Sun et al., 2022) transform point clouds into sequences, most notably PointTransformer v3 (PTv3) (Wu et al., 2024b), which projects points into voxels and orders them using space-filling curves (e.g., Hilbert curve). While scalable, these curves introduce artificial discontinuities that can break local spatial relationships.
@@
Rather than using a hierarchical decomposition, another line of work proposes cluster attention (Janny et al., 2023; Alkin et al., 2024a; Wu et al., 2024a). These methods first group points into clusters and aggregate their features at the cluster centroids through message passing or cross-attention. After computing attention between the centroids, the updated features are then distributed back to the original points. While these approaches yield the quadratic cost only in the number of clusters, they introduce an information bottleneck at the clustering step that may sacrifice fine-grained details and fail to capture features at multiple scales - a limitation our hierarchical approach aims to overcome.
Beyond attention, several alternatives have been developed for processing large-scale systems with irregular geometries. Message-passing neural networks (Gilmer et al., 2017) typically address scalability through multi-level graph representations (Lam et al., 2023; Cao et al., 2023; Valencia et al., 2025). Additionally, recent works have explored sub-quadratic convolution-based architectures like Hyena (Moskalev et al., 2025) and state-space models (Zhang et al., 2025), which offer promising alternatives for efficient processing of geometric data without the computational overhead of attention mechanisms.
@@ 3.2 Ball tree
To enable efficient implementation, we want to work with perfect binary trees, i.e., trees where all internal nodes have exactly two children and all leaf nodes appear at the same depth. To achieve this, we pad the leaf level of a ball tree with virtual nodes, yielding the total number of nodes $2^{m}$, where $m=\text{ceil}(\log_{2}(n))$.
@@ 3.2.1 Ball tree properties
The contiguous storage property, combined with the fixed size of balls at each level, enables efficient implementation through tensor operations. Specifically, accessing any ball $B\in L_{i}$ simply requires selecting a contiguous sequence of $2^{i}$ indices. For instance, in the example above, for $i=2$, we select a:d and e:h to access the balls. Since the balls are equal in size, we can simply reshape $L_{0}$ to access any level. This representation makes it particularly efficient to implement our framework’s core operations - ball attention and coarsening/refinement - which we will introduce next.
Another important property of ball trees is that while they cover the whole point set, they are not required to partition the entire space. Coupled with completeness, it means that at each tree level, the nodes are essentially associated with the same scale. This contrasts with other structures such as octrees that cover the entire space and whose nodes at the same level can be associated with regions of different sizes:
@@
Figure 2: Ball tree vs. octree construction. Colors highlight the difference in scales for nodes including the same number of points.
@@ 4.1 Ball tree attention
For each ball attention operator, we specify a level $k$ of the ball tree where each ball $B\in L_{k}$ contains $2^{k}$ leaf nodes. The choice of $k$ presents a trade-off: larger balls capture longer-range dependencies, while smaller balls are more resource-efficient. For each ball $B\in L_{k}$, we collect the leaf nodes within B:
@@
along with their features from $X_{\mathrm{leaf}}$:
@@
As attention is computed independently for each ball $B\in L_{k}$, the computational cost is reduced from quadratic to linear. Precisely, for ball attention, the complexity is $\mathcal{O}(|B|^{2}\cdot\frac{n}{|B|})$, i.e., quadratic in the ball size and linear in the number of balls:
@@
with a learnable parameter $\sigma\in\mathbb{R}$ (Wessels et al., 2024). The term decays rapidly as the distance between two nodes increases, which enforces locality and helps to mitigate potential artifacts from the tree building, particularly in cases where distant points are grouped together.
@@
To increase the receptive field of our attention operator, we implement cross-ball connections inspired by the shifted window approach in Swin Transformer (Liu et al., 2021). There, patches are displaced diagonally by half their size to obtain two different image partitioning configurations. This operation can be equivalently interpreted as keeping the patches fixed while sliding the image itself.
Following this interpretation, we rotate the point cloud and construct the second ball tree $T_{\mathrm{rot}}=\{L_{0}^{\mathrm{rot}},...,L_{m}^{\mathrm{rot}}\}$, which induces a permutation $\pi^{\mathrm{rot}}$ of leaf nodes (see Fig. 3, center). We can then compute ball attention on the rotated configuration by first permuting the features according to $\pi^{\mathrm{rot}}$, applying attention, and then permuting back:
@@
The inverse operation, refinement, allocates information from a coarse representation back to finer scales. More precisely, for a ball $B\in L_{k}$, its features are distributed back to the nodes at level $L_{k-l}$ contained within $B$ as:
@@ 4.2 Model architecture
We are now ready to describe the details of the main model to which we refer as Erwin [^4] We pay homage to Swin Transformer as our model is based on rotating windows instead of sliding, hence Rwin $\rightarrow$ Erwin. (see Fig. 3) - a hierarchical transformer operating on ball trees.
@@
At the embedding phase, we first construct a ball tree on top of the input point cloud and pad the leaf layer to complete the tree, as described in Section 3.2. To capture local geometric features, we employ a small-scale MPNN, which is conceptually similar to PointTransformer’s embedding module using sparse convolution. When input connectivity is not provided (e.g., mesh), we utilize the ball tree structure for a fast nearest neighbor search.
@@
Following a UNet structure (Ronneberger et al., 2015; Wu et al., 2024b), Erwin processes features at multiple scales through encoder and decoder paths (Fig. 3, right). The encoder progressively coarsens the ball tree while increasing feature dimensionality to maintain expressivity. The coarsening factor is a hyperparameter that takes values that are powers of $2$. At the decoder stage, the representation is refined back to the original resolution, with skip connections from corresponding encoder levels enabling multi-scale feature integration.
@@ 5 Experiments
The code is available at https://github.com/maxxxzdn/erwin. We summarize datasets in Table 1 and provide implementation details in Appendix B.
Table 1: Summary of benchmark datasets. $\dagger$ indicates varying size, for which we report the average number of nodes.
Geometry Benchmarks #Dim #Nodes
Point Cloud Elasticity 2D 972
Cosmology 3D 5,000
MD 3D+Time 890
Structured Plasticity 2D+Time 3,131
Airfoil 2D 11,271
Pipe 2D 16,641
Unstructured Shape-Net Car 3D 32,186
EAGLE 2D+Time 3,388
@@
To experimentally evaluate Erwin’s scaling, we learn the power-law [^5] We only use data for $n\geq 1024$ to exclude overhead costs. form $\mathrm{Runtime}=C\cdot n^{\beta}$ by first applying the logarithm transform to both sides and then using the least squares method to evaluate $\beta$. The result is an approximately linear scaling with $\beta$$=$$1.054$ with $R^{2}$$=$$0.999$, see Fig. 5. Ball tree construction accounts for $<$$5$% of the overall time (see Tables 7,8), proving the efficiency of our method for linearizing attention for point clouds.
@@
One of the theoretical properties of our model is that with sufficiently many layers, its receptive field is global. To verify this claim experimentally, for an arbitrary target node, we run the forward pass of Erwin and MPNN and compute gradients of the node output with respect to all input nodes’ features. If the gradient is non-zero, the node is considered to be in the receptive field of the target node. The visualization is provided in Fig. 5, right, where we compare the receptive field of our model with that of MPNN. As expected, the MPNN has a limited receptive field, as it cannot exceed $N$ hops, where $N$ is the number of message-passing layers. Conversely, Erwin implicitly computes all-to-all interactions, enabling it to capture long-range interactions in the data.
@@ 5.1 Cosmological simulations
To demonstrate our model’s ability to capture long-range interactions, we use the cosmology benchmark (Balla et al., 2024), which consists of large-scale point clouds representing potential galaxy distributions.
@@
The dataset is derived from N-body simulations that evolve dark matter particles from the early universe to the present time. After the simulation, gravitationally bound structures (halos) are identified, from which the 5,000 heaviest ones are selected as potential galaxy locations. The halos form local clusters through gravity while maintaining long-range correlations that originated from interactions in the early universe, reflecting the initial conditions.
@@
The input is a point cloud $\mathbf{X}\in\mathbb{R}^{5000\times 3}$, where each row corresponds to a galaxy and each column to $x,y,z$ coordinate respectively. The task is a regression problem to predict the velocity of every galaxy $\mathbf{Y}\in\mathbb{R}^{5000\times 3}$. We vary the size of the training dataset from 64 to 8,192, while the validation and test datasets have a fixed size of 512. The models are trained using the mean squared error between predicted and ground truth velocities.
@@ Results
The results are shown in Fig. 6. We compare against multiple equivariant (NequIP (Batzner et al., 2021), SEGNN (Brandstetter et al., 2022)) and non-equivariant (MPNN (Gilmer et al., 2017), PTv3 (Wu et al., 2024b)) baselines. In the small data regime, graph-based equivariant models are preferable. However, as the training set size increases, their performance plateaus. We note that this is also the case for non-equivariant MPNNs, suggesting the issue might arise from failing to capture medium to large-scale interactions, where increased local expressivity has minimal impact. Conversely, transformer-based models scale favorably with training set size and eventually surpass graph-based models, highlighting their ability to capture both small and large-scale interactions. Our model demonstrates particularly strong performance and significantly outperforms other baselines for larger training sets.